OpenStudy (chycora):

When Elyse moved last weekend, the circular glass top for an end table broke. To replace it, she bought a circular glass top with a diameter of 16 inches. What is the approximate area of the circular glass top Elyse bought? (Use 3.14 for .) 803.84 in2 100.48 in2 200.96 in2 50.24 in2

OpenStudy (anonymous):

2R = D where 'd' is given 2r = 16 r = 16/2 r = 8inches

OpenStudy (chycora):

here is the image https://app158.studyisland.com/pics/185178num9.png

OpenStudy (anonymous):

area of a circular body is \[A = 2*\pi*r^2\]

OpenStudy (chycora):

ok so wat find all of the answers or wat lol

OpenStudy (anonymous):

I am only finding the area.

OpenStudy (anonymous):

\[A = 2 \times 3.14 \times 8 \times 8\]

OpenStudy (anonymous):

\[A = 6.28 \times 64\]

OpenStudy (chycora):

o ok thanx luv

OpenStudy (anonymous):

Medal please :P

OpenStudy (chycora):

so A=6.28×64=401.92 @KenichiIvy

OpenStudy (anonymous):

Yes!

OpenStudy (chycora):

ok but wat kn cuz that's nt the answer